OpenStudy (anonymous):

Two supplementary angles are in the ratio of 1 to 3. The measure of the larger angle is 60 67.5 120 135

Directrix (directrix):

By definition, supplementary angles are angles with measures that sum to 180. Because the angles are in a one to three ratio, let one angle be x and the other be 3x. x/3x = 1/3 which is one to three so the ratio has be preserved.

Directrix (directrix):

Solve this equation for x: x + 3x = 180. Remember that you are to choose the larger of the two supplementary angles as your answer. @EmilyRose373